# Amplitude Expansion Phase Field Crystal (APFC) Modeling based Efficient Dislocation Simulations using Fourier Pseudospectral Method

### v2

@@ 2 APFC Model
    $\frac{\delta f^{s}_{\rm bcc}}{\delta\eta_{i}^{*}}$ $=-2\tau(\eta_{k}\eta_{n}^{*}+\eta_{j}\eta_{l})+6v(\eta_{k}\eta_{l}\eta_{m}+\eta_{j}\eta_{m}^{*}\eta_{n}^{*}),$
    $\frac{\delta f^{s}_{\rm bcc}}{\delta\eta_{l}^{*}}$ $=-2\tau(\eta_{m}^{*}\eta_{n}^{*}+\eta_{i}\eta_{j}^{*})+6v(\eta_{i}\eta_{k}^{*}\eta_{m}^{*}+\eta_{k}\eta_{j}^{*}\eta_{n}^{*}),$
@@ 3 Numerical Method for the APFC Model
Given the lattice symmetry and periodic boundary conditions, the Fourier pseudospectral method in the spatial direction is well-suited for nonlinear and high-order PDEs due to its high accuracy and straightforward implementation. In Section 3.1, we introduce the basic setup for the Fourier pseudospectral method. Then, in Section 3.2, we provide a detailed explanation of its principles and practical implementation for solving the APFC model (2.13) using Fourier pseudospectral method.
@@ 3.1 Basic Setup
In pseudospectral methods, PDEs are solved at collocation points in physical space, where spatial derivatives are computed using orthogonal functions like trigonometric and Chebyshev polynomials. These functions are evaluated using techniques such as matrix multiplication, fast Fourier transform (FFT), or convolution. The Fourier pseudospectral method used in this study follows the procedure outlined in [45], which include selecting collocation points, discretizing the test function by computing the inner product of the equation, obtaining Fourier coefficients, and performing an inverse Fourier transform to obtain the desired function [46].
@@
$K_{N}^{d}:=\{\bm{k}=(k_{j})_{j=1}^{d}\in\mathbb{Z}^{d}:-\frac{N}{2}\leq k_{j}<\frac{N}{2}\},$
where $d$ represents dimension. Let $\mathbb{T}^{d}:=(\mathbb{R}/2\pi\mathbb{Z})^{d}$ be the $d$-dimensional torus. We discretize $\mathbb{T}^{d}$ using the fundamental domain $[0,2\pi)^{d}$ with uniform spatial discretization parameter $h=2\pi/N$. The grid $\mathbb{T}^{d}_{N}$ consists of points $\bm{r_{p}}=(x_{1,l_{1}},x_{2,l_{2}},\ldots,x_{d,l_{d}})$, where $x_{i,l_{i}}=l_{i}h$, for $0\leq l_{i}<N$, $l_{i}\in\mathbb{N}_{0}$.
@@
$\mathcal{G}_{N}:=\big{\{}u:\mathbb{T}_{N}^{d}\mapsto\mathbb{C}~{}|~{}u\text{ is }\mathbb{T}^{d}_{N}-\text{periodic}\big{\}}.$
@@
$\langle e^{\imath\bm{k}\cdot\bm{r_{p}}},e^{\imath\bm{l}\cdot\bm{r_{p}}}\rangle_{N}=\sum_{\bm{r_{p}}\in\mathbb{T}^{d}_{N}}e^{\imath(\bm{k}-\bm{l})\cdot\bm{r_{p}}}=\left\{\begin{aligned} 1,\quad&\bm{k}=\bm{l}+N\bm{m},\text{ }\bm{m}\in\mathbb{Z}^{d}\\ 0,\quad&{\rm otherwise}\\ \end{aligned}\right..$ (3.1)
@@
$\widehat{u}_{\bm{k}}=\langle u,e^{\imath\bm{k}\cdot\bm{r_{p}}}\rangle_{N},\quad\bm{k}\in K_{N}^{d}.$
We can also perform the discrete Fourier transform of periodic function $u$ by
$u(\bm{r_{p}})=\sum_{\bm{k}\in K^{d}_{N}}\widehat{u}_{\bm{k}}e^{\imath\bm{k}\cdot\bm{r_{p}}},\quad\bm{r_{p}}\in\mathbb{T}_{N}^{d}.$ (3.2)
The trigonometric interpolation of periodic function is
$I_{N}u(\bm{r})=\sum_{\bm{k}\in K^{d}_{N}}\widehat{u}_{\bm{k}}e^{\imath\bm{k}\cdot\bm{r}}.$
It is straightforward to see that $I_{N}u(\bm{r_{p}})=u(\bm{r_{p}})$. In practice, one can use the $d$-dimensional FFT to obtain the Fourier coefficients $\widehat{u}_{\bm{k}}$ in (3.2).
Remark 3.1.
In addition to the Fourier spectral method discussed above, the finite element method is also commonly used for solving the APFC model [30, 24]. However, it demands more computational resources and requires careful consideration of spatial discretization. For the sake of completeness, we include the implementation details of finite element method in solving the APFC model in Appendix C.
@@ 3.2 Implementation of Fourier pseudospectral method
In this section, we give a detailed implementation of solving the APFC model (2.13) by employing Fourier pseudospectral method.
To that end, we first expand the amplitude functions $\eta_{j}$ to
$\eta_{j}(\bm{r})=\phi_{j}e^{\imath\bm{k_{j}}\cdot\bm{u}(\bm{r})},\quad j=1,\ldots,N,$ (3.3)
@@
where $\bm{u}(\bm{r})$ corresponds to the displacement field with respect to the reference lattice [24]. The equations of each $\eta_{j}$ are complex functions, and they are independent of each other. Substituting (3.3) into (2.2), one can obtain
$n(\bm{r})$ $=n_{0}(\bm{r})+\sum_{j=1}^{N}\big{[}\phi_{j}e^{\imath\bm{k_{j}}\cdot\bm{r}}e^{\imath\bm{k_{j}}\cdot\bm{u(r)}}+\phi_{j}e^{-\imath\bm{k_{j}}\cdot\bm{r}}e^{-\imath\bm{k_{j}}\cdot\bm{u(r)}}\big{]}$ (3.4)
$=n_{0}(\bm{r})+\sum_{j=1}^{N}\big{[}\phi_{j}e^{\imath\bm{k_{j}}\cdot(\bm{r}+\bm{u(r)})}+\phi_{j}e^{-\imath\bm{k_{j}}\cdot(\bm{r}+\bm{u(r)})}\big{]}.$ (3.5)
@@
In comparison with the density defined by (2.2), we observe that $u(\bm{r})$ represents the perturbation term, while $\phi_{j}e^{\imath\bm{k_{j}}\cdot\bm{r}}$ corresponds to the reference lattice. Here, $\phi_{j}$ represents the real value of the amplitude in the relaxed crystal. These quantities can be computed by minimizing the energy functional in (2.1), which is the equilibrium lattice state of the system.
@@
In our work, square and cube are employed as computational domains for two-dimensional and three-dimensional cases, respectively. Let $L$ be the side length of computational domain. The grid points are $\bm{r_{p}}=(\frac{m_{1}L}{N},\frac{m_{2}L}{N})\in\mathbb{T}_{N}^{2}$, $0\leq m_{1},m_{2}<N$, and $\bm{r_{p}}=(\frac{m_{1}L}{N},\frac{m_{2}L}{N},\frac{m_{3}L}{N})\in\mathbb{T}_{N}^{3}$, $0\leq m_{1},m_{2},m_{3}<N$, $m_{i}\in\mathbb{N}_{0}$. The dimensions of $K_{N}^{n}$ in different situations are
$\dim(K_{N}^{n})=N^{d}=:D,\quad d=2,3.$ (3.6)
@@
The amplitude functions $\eta_{j}$ are periodic on the Bravais lattice. Based on the relationship between the Bravais lattice and the reciprocal lattice, as well as their mathematical descriptions provided in Appendix A, we can derive their discrete Fourier transform from (A.5) as follows,
$\eta_{j}(\bm{r_{p}},t)\approx I_{N}^{n}\eta_{j}(\bm{r_{p}},t)=\sum_{\bm{k}\in K_{N}^{n}}\widehat{\eta_{j}}_{\bm{k}}(t)e^{\imath(B\bm{k})\cdot\bm{r_{p}}}\quad p=0,1,\ldots,N^{d}-1,$
@@
where $\widehat{\eta_{j}}_{\bm{k}}(t)=\langle\eta_{j},e^{\imath\bm{k}\cdot\bm{r_{p}}}\rangle_{N}.$ The first and second order derivative of $\eta_{j}$ in Fourier pesduo-spectral method is defined as
$\nabla\eta_{j}\approx\sum_{\bm{k}\in K_{N}^{n}}(B\bm{k})\widehat{\eta_{j}}_{\bm{k}}(t)e^{\imath(B\bm{k})\cdot\bm{r_{p}}},$
$\nabla^{2}\eta_{j}\approx-\sum_{\bm{k}\in K_{N}^{n}}|B\bm{k}|^{2}\widehat{\eta_{j}}_{\bm{k}}(t)e^{\imath(B\bm{k})\cdot\bm{r_{p}}}.$
@@
Substituting them into the APFC model (2.13), we have
$\sum_{\bm{k}\in K_{N}^{n}}\frac{\partial\widehat{\eta_{j}}_{\bm{k}}(t)}{\partial t}e^{\imath(B\bm{k})\cdot r_{p}}=$ $-|\bm{k_{j}}|^{2}\left[\sum_{\bm{k}\in K_{N}^{n}}\Big{(}B_{0}+B_{x}\big{(}-(B\bm{k})^{2}-2\bm{k_{j}}\cdot({B\bm{k}})\big{)}^{2}\Big{)}\widehat{\eta_{j}}_{\bm{k}}(t)e^{\imath(B\bm{k})\cdot r_{p}}+\widehat{G}(\eta)\right],$
@@
$G(\eta)=G(t,\eta_{j}(\bm{r_{p}},t)):=3v(A^{2}-|\eta_{j}|^{2})\eta_{j}+\frac{\delta f^{s}}{\delta\eta_{j}^{*}}$
represents the function related to $\eta_{1},\ldots,\eta_{N},\eta_{1}^{*},\ldots,\eta_{N}^{*}$ and is treated explicitly in the time discretization. By using the basis function $e^{\imath(Bk)\cdot r_{q}}$ to perform the inner product at both ends of the equation simultaneously and applying orthogonality (3.1) yields
$\frac{\partial\widehat{\eta}_{j_{\bm{k}}}}{\partial t}=-|\bm{k_{j}}|^{2}\big{[}B_{0}+B_{x}\big{(}-(B\bm{k})^{2}-2\bm{k_{j}}\cdot({B\bm{k})}\big{)}^{2}\big{]}\widehat{\eta}_{j_{\bm{k}}}(t)+\widehat{G}_{\bm{k}}(\eta).$ (3.7)
A semi-implicit scheme is employed for time discretization, allowing us to obtain,
$\frac{\widehat{\eta}_{j_{\bm{k}}}^{(n+1)}-\widehat{\eta}_{j_{\bm{k}}}^{(n)}}{\Delta t}=-|\bm{k_{j}}|^{2}\big{[}B_{0}+B_{x}\big{(}-(B\bm{k})^{2}-2\bm{k_{j}}\cdot({B\bm{k})}\big{)}^{2}\big{]}\widehat{\eta}_{j_{\bm{k}}}^{n+1}+\widehat{G}_{\bm{k}}(\eta^{(n)}).$
@@
where $\Delta t$ is the time step size. The iterative format is then given by
$\widehat{\eta}_{j_{\bm{k}}}^{(n+1)}=\frac{\widehat{\eta}_{j_{\bm{k}}}^{(n)}-\Delta t|\bm{k_{j}}|^{2}\widehat{G}_{\bm{k}}(\eta^{(n)})}{1+\Delta t|\bm{k_{j}}|^{2}\big{[}B_{0}+B_{x}\big{(}-(B\bm{k})^{2}-2\bm{k_{j}}\cdot({B\bm{k})}\big{)}^{2}\big{]}}.$ (3.8)
@@
Next, we analyze the computational complexity of solving (3.8) at each time step. The iterative format can be efficiently computed using the Fourier pseudospectral method in $d$-dimensional reciprocal space. While linear terms are straightforward, the nonlinear terms $\widehat{G}(\eta)$ in (3.7) involve computationally intensive convolutions. However, these can be efficiently calculated as dot products in physical space using the Fast Fourier Transform (FFT) [43]. Thus, the complexity of the Fourier pseudospectral method for solving (3.8) is reduced to $\mathcal{O}(D\log D)$ per time step, where $D$ is the number of degrees of freedom defined in (3.6). This method achieves high numerical efficiency and accelerates computations, as will be demonstrated in the following section.

### v3

@@ 2 APFC Model
    $\frac{\delta f^{s}_{\rm bcc}}{\delta\eta_{i}^{*}}$ $=-2\gamma(\eta_{k}\eta_{n}^{*}+\eta_{j}\eta_{l})+6v(\eta_{k}\eta_{l}\eta_{m}+\eta_{j}\eta_{m}^{*}\eta_{n}^{*}),$
    $\frac{\delta f^{s}_{\rm bcc}}{\delta\eta_{l}^{*}}$ $=-2\gamma(\eta_{m}^{*}\eta_{n}^{*}+\eta_{i}\eta_{j}^{*})+6v(\eta_{i}\eta_{k}^{*}\eta_{m}^{*}+\eta_{k}\eta_{j}^{*}\eta_{n}^{*}),$
@@ 3 Numerical Method for the APFC Model
Given the lattice symmetry and periodic boundary conditions, the Fourier pseudospectral method in the spatial direction is well-suited for nonlinear and high-order PDEs due to its high accuracy and straightforward implementation. In Section 3.1, we introduce the basic setup for the Fourier pseudospectral method. Then, in Section 3.2, we provide a detailed explanation of its principles and practical implementation for solving the APFC model (13) using the Fourier pseudospectral method.
@@ 3.1 Basic setup
In pseudospectral methods, PDEs are solved at collocation points in physical space, where spatial derivatives are computed using orthogonal functions like trigonometric and Chebyshev polynomials. These functions are evaluated using techniques such as matrix multiplication, fast Fourier transform (FFT), or convolution. The Fourier pseudospectral method used in this study follows the procedure outlined in jiang2024numerical , which includes selecting collocation points, discretizing the test function by computing the inner product of the equation, obtaining Fourier coefficients, and performing an inverse Fourier transform to obtain the desired function 2011Spectral .
@@
$K_{N}^{d}:=\left\{\bm{k}=(k_{j})_{j=1}^{d}\in\mathbb{Z}^{d}:-\frac{N}{2}\leq k_{j}<\frac{N}{2}\right\},$
where $d$ represents dimension. Let $\mathbb{T}^{d}:=(\mathbb{R}/2\pi\mathbb{Z})^{d}$ be the $d$-dimensional torus. We discretize $\mathbb{T}^{d}$ using the fundamental domain $0,2\pi)^{d}$ with uniform spatial discretization parameter $h=2\pi/N$. The grid $\mathbb{T}^{d}_{N}$ consists of points $\bm{r}_{p}=(x_{1,l_{1}},x_{2,l_{2}},\ldots,x_{d,l_{d}})$, where $x_{i,l_{i}}=l_{i}h$, for $0\leq l_{i}<N$, $l_{i}\in\mathbb{N}_{0}$.
@@
$\mathcal{G}_{N}:=\big\{u:\mathbb{T}_{N}^{d}\mapsto\mathbb{C}~\lvert~u\text{ is }\mathbb{T}^{d}_{N}-\text{periodic}\big\}.$
@@
$\langle e^{\imath\bm{k}\cdot\bm{r}_{p}},e^{\imath\bm{l}\cdot\bm{r}_{p}}\rangle_{N}=\sum_{\bm{r}_{p}\in\mathbb{T}^{d}_{N}}e^{\imath(\bm{k}-\bm{l})\cdot\bm{r}_{p}}=\left\{\begin{aligned} 1,\quad&\bm{k}=\bm{l}+N\bm{m},\text{ }\bm{m}\in\mathbb{Z}^{d}\\ 0,\quad&{\rm otherwise}\\ \end{aligned}\right.$. (17)
@@
$u_{\bm{k}}=\langle u,e^{\imath\bm{k}\cdot\bm{r}_{p}}\rangle_{N},\quad\bm{k}\in K_{N}^{d}.$ (18)
We can also perform the discrete Fourier transform of the periodic function $u$ by
$u(\bm{r}_{p})=\sum_{\bm{k}\in K^{d}_{N}}u_{\bm{k}}e^{\imath\bm{k}\cdot\bm{r}_{p}},\quad\bm{r}_{p}\in\mathbb{T}_{N}^{d}.$ (19)
The trigonometric interpolation of the periodic function is
$I_{N}u(\bm{r})=\sum_{\bm{k}\in K^{d}_{N}}u_{\bm{k}}e^{\imath\bm{k}\cdot\bm{r}}.$ (20)
It is straightforward to see that $I_{N}u(\bm{r}_{p})=u(\bm{r}_{p})$. In practice, one can use the $d$-dimensional FFT to obtain the Fourier coefficients $u_{\bm{k}}$ in (19).
Remark 4
In addition to the Fourier spectral method discussed above, the finite element method is also commonly used for solving the APFC model 2017Controlling . However, it demands more computational resources and requires careful consideration of spatial discretization. For the sake of completeness, we include the implementation details of finite element method in solving the APFC model in Appendix C.
@@ 3.2 Implementation of Fourier pseudospectral method
In this section, we present a detailed implementation of solving the APFC model (13) by employing Fourier pseudospectral method.
We begin by expanding the amplitude functions $\eta_{j}$ to
$\eta_{j}(\bm{r},t)=\phi_{j}e^{\imath\bm{k}_{j}\cdot\bm{u}(\bm{r})},\quad j=1,\ldots,M,$ (21)
@@
where $\bm{u}(\bm{r})$ corresponds to the displacement field with respect to the reference lattice 2017Controlling . The equations of each $\eta_{j}$ are complex functions, and they are independent of each other. Substituting (21) into (2), we can obtain
$n(\bm{r},t)$ $=n_{0}(\bm{r},t)+\sum_{j=1}^{M}\big[\phi_{j}e^{\imath\bm{k}_{j}\cdot\bm{r}}e^{\imath\bm{k}_{j}\cdot\bm{u}(\bm{r})}+\phi_{j}e^{-\imath\bm{k}_{j}\cdot\bm{r}}e^{-\imath\bm{k}_{j}\cdot\bm{u}(\bm{r})}\big]$ (22)
$=n_{0}(\bm{r},t)+\sum_{j=1}^{M}\big[\phi_{j}e^{\imath\bm{k}_{j}\cdot(\bm{r}+\bm{u}(\bm{r}))}+\phi_{j}e^{-\imath\bm{k}_{j}\cdot(\bm{r}+\bm{u}(\bm{r}))}\big].$ (23)
@@
Compared with the density defined by (2), we observe that $\bm{u}(\bm{r})$ represents the perturbation term, while $\phi_{j}e^{\imath\bm{k}_{j}\cdot\bm{r}}$ corresponds to the reference lattice. Here, $\phi_{j}$ represents the real value of the amplitude in the relaxed crystal. These quantities can be computed by minimizing the energy functional in (1), which is the equilibrium lattice state of the system.
@@
In our work, square and cube are employed as computational domains for two-dimensional and three-dimensional cases, respectively. Let $L$ be the side length of the computational domain. The grid points are $\bm{r}_{p}=(\frac{m_{1}L}{N},\frac{m_{2}L}{N})\in\mathbb{T}_{N}^{2}$, $0\leq m_{1},m_{2}<N$, and $\bm{r}_{p}=(\frac{m_{1}L}{N},\frac{m_{2}L}{N},\frac{m_{3}L}{N})\in\mathbb{T}_{N}^{3}$, $0\leq m_{1},m_{2},m_{3}<N$, $m_{i}\in\mathbb{N}_{0}$. The dimensions of $K_{N}^{d}$ in different situations are
$\dim(K_{N}^{d})=N^{d}=:D,\quad d=2,3.$ (24)
@@
The amplitude functions $\eta_{j}$ are periodic on the Bravais lattice. Based on the relationship between the Bravais lattice and the reciprocal lattice, as well as their mathematical descriptions provided in Appendix A, we can derive their discrete Fourier transform from (49) as follows:
$\eta_{j}(\bm{r}_{p},t)\approx I_{N}^{d}\eta_{j}(\bm{r}_{p},t)=\sum_{\bm{k}\in K_{N}^{d}}\eta_{j_{\bm{k}}}(t)e^{\imath(B\bm{k})\cdot\bm{r}_{p}},\quad p=0,1,\ldots,N^{d}-1,$
@@
where $\eta_{j_{\bm{k}}}(t)=\langle\eta_{j},e^{\imath\bm{k}\cdot\bm{r}_{p}}\rangle_{N}.$ The first-order and second-order derivatives of $\eta_{j}$ in the Fourier pseudospectral method are given by
$\nabla\eta_{j}$ $\approx\sum_{\bm{k}\in K_{N}^{d}}(B\bm{k})\eta_{j_{\bm{k}}}(t)e^{\imath(B\bm{k})\cdot\bm{r}_{p}},$ (25)
$\nabla^{2}\eta_{j}$ $\approx-\sum_{\bm{k}\in K_{N}^{d}}\lvert B\bm{k}\rvert^{2}\eta_{j_{\bm{k}}}(t)e^{\imath(B\bm{k})\cdot\bm{r}_{p}}.$
3.2.1 Time discretization
@@
We now proceed to the discretization of the APFC dynamics. Substituting (25) into the APFC model (13) yields
$\sum_{\bm{k}\in K_{N}^{d}}\frac{\partial\eta_{j_{\bm{k}}}(t)}{\partial t}e^{\imath(B\bm{k})\cdot\bm{r}_{p}}$
$=$ $-\lvert\bm{k}_{j}\rvert^{2}\left[\sum_{\bm{k}\in K_{N}^{d}}\Big(B_{0}+B_{x}\big(-(B\bm{k})^{2}-2\bm{k}_{j}\cdot({B\bm{k}})\big)^{2}\Big)\eta_{j_{\bm{k}}}(t)e^{\imath(B\bm{k})\cdot\bm{r}_{p}}+G_{\bm{k}}(\eta)\right],$
@@
$G(\eta)=G\big(t,\eta_{j}(\bm{r}_{p},t)\big):=3v(A^{2}-\lvert\eta_{j}\rvert^{2})\eta_{j}+\frac{\delta f^{s}}{\delta\eta_{j}^{*}}$
denotes the function depending on $\eta_{1},\ldots,\eta_{M},\eta_{1}^{*},\ldots,\eta_{M}^{*}$; this function is treated explicitly in the time discretization.
Taking the inner product on both sides of the equation with the basis function $e^{\imath(B\bm{k})\cdot\bm{r}_{q}}$ and applying the orthogonality relation (17), we obtain
$\frac{\partial\eta_{j_{\bm{k}}}}{\partial t}=-\lvert\bm{k}_{j}\rvert^{2}\big[B_{0}+B_{x}\big(-(B\bm{k})^{2}-2\bm{k}_{j}\cdot(B\bm{k})\big)^{2}\big]\eta_{j_{\bm{k}}}(t)+G_{\bm{k}}(\eta).$ (26)
For the time discretization, we employ a semi-implicit scheme, leading to
$\frac{\tilde{\eta}_{j_{\bm{k}}}^{(n+1)}-\tilde{\eta}_{j_{\bm{k}}}^{(n)}}{\Delta t}=-\lvert\bm{k}_{j}\rvert^{2}\big[B_{0}+B_{x}\big(-(B\bm{k})^{2}-2\bm{k}_{j}\cdot(B\bm{k})\big)^{2}\big]\tilde{\eta}_{j_{\bm{k}}}^{(n+1)}+\tilde{G}_{\bm{k}}(\eta^{(n)}),$
@@
where $\Delta t$ is the time step size, and $\tilde{\eta}_{j_{\bm{k}}}^{(n)}=\tilde{\eta}_{j_{\bm{k}}}(t_{n})$ is the numerical solution at time $t_{n}=n\Delta t$. The resulting iterative scheme is given by
$\tilde{\eta}_{j_{\bm{k}}}^{(n+1)}=\frac{\tilde{\eta}_{j_{\bm{k}}}^{(n)}-\Delta t\lvert\bm{k}_{j}\rvert^{2}\tilde{G}_{\bm{k}}(\eta^{(n)})}{1+\Delta t\lvert\bm{k}_{j}\rvert^{2}\big[B_{0}+B_{x}\big(-(B\bm{k})^{2}-2\bm{k}_{j}\cdot(B\bm{k})\big)^{2}\big]}.$ (27)
The approximate solution $\tilde{\eta}_{j}(\bm{r},t_{n})$ is then reconstructed via
$\tilde{\eta}_{j}(\bm{r},t_{n})=\sum_{\bm{k}\in K_{N}^{d}}\tilde{\eta}_{j_{\bm{k}}}(t_{n})e^{\imath(B\bm{k})\cdot\bm{r}}.$ (28)
The procedure for calculating the displacement field $\bm{u}(\bm{r})$ from the numerical solutions $\tilde{\eta}_{j}$ is presented in Sections 4.1.2 and 4.2.2.
Given the analyticity of $\eta_{j}$ established in Assumption 1 (see the discussion in Remark 3), the Fourier pseudospectral method achieves spectral accuracy 2011Spectral . We now present the corresponding interpolation error estimate in the $L^{2}$-norm.
Theorem 3.1
Under the analyticity Assumption 1, let $T>0$ be sufficiently large and $\eta_{j}$ denote the solution of (13). There exist positive constants $C_{\rm s}$ and $c$, independent of $\eta_{j}$ and $N$, such that
$\|I_{N}\eta_{j}(\cdot,T)-\eta_{j}(\cdot,T)\|_{L^{2}}\leq C_{\rm s}e^{-cN},$
where $I_{N}$ is the interpolation operator defined in (20).
Remark 5
To rigorously establish the convergence of the fully discrete numerical solution $\tilde{\eta}_{j}(\cdot,t_{n})$ to the exact solution $\eta_{j}(\cdot,t_{n})$, a comprehensive error estimate that accounts for both spatial and temporal discretizations is required. Such a complete analysis typically hinges on a detailed investigation of the scheme’s energy dissipation properties. While our numerical experiments in Section 4 clearly demonstrate exponential decay over long time intervals, a full mathematical justification of this behavior lies beyond the scope of the present paper and is therefore deferred to future work.
@@
We now analyze the computational complexity of solving the iterative scheme (27) at each time step. This formulation can be efficiently implemented using the Fourier pseudospectral method in $d$-dimensional reciprocal space. While the linear terms are handled directly, the nonlinear contributions $G_{\bm{k}}(\eta)$ in (26) involve convolution operations that would be computationally expensive if evaluated directly in Fourier space. However, these nonlinear terms can be efficiently computed as local dot products in physical space by leveraging the Fast Fourier Transform (FFT) algorithm zhang2008efficient . Consequently, the overall complexity per time step is reduced to $\mathcal{O}(D\log D)$, where $D$ denotes the number of degrees of freedom defined in (24). This approach achieves high numerical efficiency and significantly accelerates computations, as will be demonstrated in the following section.